OpenStudy (anonymous):

Choose exactly two answers that are correct. 3. Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 10 + (6 / 3) = 10 + 2 = 12 B. 16 - (4 + 7) = 12 + 7 = 19 C. 7* 6 + 4 = 7 * 10 = 70 D. 2 + 5 * 7 = 2 + 35 = 37

OpenStudy (anonymous):

D & A

OpenStudy (anonymous):

work?

OpenStudy (anonymous):

D: 2+5*7=37 2+35=37 A: 10+(6/3)=12 10+2=12

OpenStudy (anonymous):

Did that help?